I have to complete the square and dont understand how you do it. Can you explain???
Ex1: X squared - 2X-5=0
EX2:Ysquared + 8y+6=0

Answer by josmiceli(6786)

Get the equation into the form (x + a)^2 -k = 0
expand
[1] x^2 + 2ax + a^2 -k = 0
[2] x^2 -2x -5 = 0
comparing coefficients of x
-2 = 2a
a = -1
a^2 = +1
substituting in [1]
x^2 -2x +1 -k = 0
In order for this = [2], k = -6
(x - 1)^2 -6 = 0
---------
y^2 +8y + 6 = 0
y^2 + 2ay + a^2 -k = 0
2a = 8
a = 4
a^2 = 16
y^2 + 8y + 16 -10 = 0
(y + 4)^2 -10 = 0